Circular No. 6289

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996A IN ANONYMOUS GALAXY B. P. Schmidt, Mount Stromlo and Siding Spring Observatories, re-

ports the discovery of a supernova by R. O. Evans on a U.K. Schmidt R film taken by P. Cass on Jan. 12 as part of the U.K. Schmidt Supernova Search. SN 1996A, of mag ∼ 18, is located at α = 5^{h}10^{m}47^{s}.1, δ = −3^{o}00′25′′ (equinox 2000.0), in an anonymous galaxy. It is offset 1′′ east and 4′′ south from a mag 14 star located on top of the galaxy’s nucleus. It is not visible on a similar film taken 1995 Dec. 12.

M. M. Phillips and A. Clocchiatti, Cerro Tololo Interamerican Observatory, report: “A fully-reduced spectrogram (range 350–720 nm, resolution 0.4 nm) of SN 1996A, obtained on Jan. 18.14 UT at the 4-m Blanco telescope, shows the typical features of a type II supernova approximately three weeks after maximum light. Hβ is blueshifted by ∼ 9000 km/s, and Hα has a well-developed emission and a weak absorption blueshifted by ∼ 9500 km/s. The parent galaxy shows a redshift of 0.033, measured from emission in Hα and [N II] 6584 nm. The supernova is offset by 9′′ in p.a. ∼ 140^{o} from the bright star located on top of the galaxy.”

COMET 67P/CHURYUMOV-GERASIMENKO Total visual magnitude estimates: 1995 Aug. 23.22 UT, 13.6 (C. S.

Morris, southern California, 0.26-m reflector); Sept. 23.22, 13.7 (Morris); Oct. 20.80, 14.2 (G. Bakos, Ráktanya, Hungary, 0.44-m reflector); 22.79, 14.3 (K. Sárneczky, Ráktanya, Hungary, 0.44-m reflector); Nov. 13.13, 13.2 (A. Hale, Cloudcroft, NM, 0.41-m reflector); 18.77, 13.1 (Sárneczky); 25.11, 12.9 (Hale); Dec. 11.09, 12.6 (Hale); 19.14, 12.5 (Hale); 21.15, 12.2 (Morris); 1996 Jan. 8.12, 11.0 (Morris); 13.18, 11.3 (Morris); 14.14, 11.3 (Morris).

COMET 58P/JACKSON-NEUJMIN Total visual magnitude estimates: 1995 Sept. 21.21 UT, 13.0 (Mor-

ris); 27.21, 13.0 (Hale); Oct. 12.12, 12.9 (Hale); 19.23, 12.5 (Hale); 22.21, 12.0 (Morris); 27.19, 12.1 (Hale); Nov. 11.11, 11.4 (Hale); 16.17, 10.6 (R. Keen, Mt. Thorodin, CO, 0.32-m reflector); 22.43, 10.6 (P. Camilleri, Cobram, Vic., Australia, 0.20-m reflector); 25.21, 10.6 (Hale); Dec. 11.11, 11.1 (Hale); 19.17, 10.9 (Hale); 22.12, 11.0 (Morris); 1996 Jan. 13.19, 12.0 (Morris); 14.14, 12.0 (Morris).

1996 January 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Brian G. Marsden